OpenStudy (kitkat1):

Medal for best answer!!!!!! MThe distance traveled, in meters, of a coin dropped from a tall building is modeled by the equation d(t) = 4.9t2 where d equals the distance traveled at time t seconds and t equals the time in seconds. What does the average rate of change of d(t) from t = 3 to t = 6 represent? (2 points) The coin travels an average distance of 44.1 meters from 3 seconds to 6 seconds. The coin falls down with an average speed of 14.7 meters per second from 3 seconds to 6 seconds. The coin falls down with an average speed of 44.1 meters per second from 3 seconds to 6 seconds.

OpenStudy (kitkat1):

The coin travels an average distance of 14.7 meters from 3 seconds to 6 seconds.

OpenStudy (kitkat1):

@rebeccaxhawaii

OpenStudy (kitkat1):

@whpalmer4

OpenStudy (kitkat1):

@foolofatook

OpenStudy (kitkat1):

I think it is either the B or D

OpenStudy (ineedhelplz):

\[\large{ \text{Average}=\cfrac{1}{6-3}\int\limits_3^6{4.9t^2dt}=\cfrac{1}{3}\times\cfrac{4.9t^3}{3}|^6_3\\ =\cfrac{4.9}{9}\times\left(6^3-3^3\right)\ }\] The average rate of change is the average velocity: \[\large{ \text{Average Velocity }=\cfrac{1}{6-3}\int\limits_3^6{\cfrac{d}{dt}4.9t^2dt}\\= \cfrac{1}{3}4.9t^2|_3^6=4.9\times\cfrac{6^2-3^2}{3}=44.1~m/s }\] So it travels at a speed of 44.1 m/s from 3 seconds to 6 seconds

OpenStudy (kitkat1):

Okay but it says to find average rate or change so 3-6=3

OpenStudy (kitkat1):

44.1 divided by 3 =14.7

OpenStudy (ineedhelplz):

Ok...

OpenStudy (kitkat1):

I just dont get from which 2 I should pick B Or D

OpenStudy (ineedhelplz):

x/t is speed. So the rate of change will be related to speed. At t=3; the coin is falling at 29.4m/s and at t-6; the coin is falling at 58.8m/s. If you take these and add them together, then divide by 2, you'll get rate per 3 seconds. Divide by 3 and you get 14.7m/s.

OpenStudy (ineedhelplz):

That's what you where saying correct?

OpenStudy (kitkat1):

Yes

OpenStudy (kitkat1):

Now what answer choice should i pick B or D?

OpenStudy (ineedhelplz):

B :)

OpenStudy (kitkat1):

Are you sure is there any difference between 2?

OpenStudy (ineedhelplz):

Yes, one says travels and the other says falls

OpenStudy (kitkat1):

Okay can you help me with one more?

OpenStudy (ineedhelplz):

Yes, but please close this question and open a new one!

OpenStudy (kitkat1):

Okay ill tag you

OpenStudy (kitkat1):

@Ineedhelplz